Mirror Cover Pneumatics Issue

Tuesday, December 15, 2015 8:46 AM

Problem(s) Encountered:

When we were closing up, one mirror cover (the one farthest from the garage door when the telescope is at zenith) did not close. We had turned the drives switch off shortly after hitting the mirror cover switch, if that might have had any effect. We solved the problem by switching the mirror cover switch to open again (which did not actually open any additional petals) waiting a bit, then flipping it back to closed again. Then the one petal finally closed itself.


In the middle of a long observation around 0500 UT/2200 local, as the dome was rotating towards Az=10 degrees (moving in decreasing Az) we heard a loud sound like a door slamming. It was not either of us. We couldn''t find anything visibly wrong with the telescope, and all the readings were normal. So we finished our exposure, shut off the auto dome, sent the

telescope to zenith and closed mirror covers (the humidity has also been dancing around, so we didn''t want to worry about that too) and inspected the dome for any problems.


The remainder of the reported issues are actually related to each other.  Checked the mirror covers this morning and as suspected, found the West cover to be the last one to have closed, meaning they were no longer in order.  I used the carabiner and rope to pull up on the west cover to allow the other petals to open.  They opened quite slowly.  The west petal opened, but seemed to have very little pressure to hold it in the “open” position.  Indeed, I could easily push it over and it would slam shut and not open back up.  This is the last petal to open in a system that effectively distributes pneumatic (air) pressure in series.  Therefore, if there is anything throttling back the air pressure upstream, this petal would show symptoms more readily than any other.  The dome was opened to allow the temperature inside to come up (it was warmer outside than in the dome) above freezing.  The pneumatic lines, as well as the cylinders, were all warmed with a hair drier.  Problem solved.  After allowing the telescope, and subsequently, the pneumatic system, to warm up a bit more through the day, the pneumatic lines from the compressor to the relays were purged, exhausting more water than I’ve ever seen from them.  This is somewhat surprising as the lines were purged yesterday as well.  Perhaps some of the condensate yesterday was in solid state?  I don’t know.  Regardless, a lot of water was removed.  The line will be purged again before we depart for the day.

In summary, I think there was a slight ice plug in the line, limiting air pressure to the cylinders, and thus limiting the holding force applied by the cylinders to keep the petals, especially the West petal, up.  I believe this also limited the closing pressure on the same petal.  I think this led to the issues exhibited when closing the telescope this morning.  In addition, I think this also led to the west petal slamming shut at some point during observations last night.  My theory from this morning was that this was the actual cause of the BANG, and not a door closing.  I held off on writing this up until I could get some more information from the observer.  The information received appears to bolster this theory—counts on a star were mysteriously lower last night than expected when running an integration, leading to the need for a longer exposure time.  All this—the BANG, the reduced flux, and the issues with closing—points to issues with pneumatic flow to the mirror cover petals resulting in one of the mirror cover petals being closed through observations.  The second BANG witnessed can also be explained by this:  the air pressure in the system was slow to come back up (~80psi).  Once it did come back up, the petal opened up again.  But with little holding pressure, fell back down as the telescope was moved to zenith.